The probabilities that Audu, Chioma and Yinka pass UTME are 1/3, 2/5 and 3/8 respectively. Find the probability that the three of them will fail the exam.
1
Expert's answer
2021-10-26T12:04:50-0400

The probabilities that Audu, Chioma and Yinka pass UTME are 1/3, 2/5 and 3/8 respectively then the probability that they all will fail the exam is:

"=(1-\frac{1}{3})\times(1-\frac{2}{5})\times(1-\frac{3}{8})"

"=(\frac{2}{3})\times(\frac{3}{5})\times(\frac{5}{8})"


"=\frac{2}{8}=\frac{1}{4}=0.25"